OpenStudy (ello1221):

You start a search for a buried object by marking the center of a field as (0, 0), with coordinates giving distances in yards. Coordinates to the north or east are positive and coordinates to the south or west are negative. You find nothing at (–12, 8), so you try a likely-looking spot 3 yards to the east and 11 yards to the south of the first spot. What are the coordinates of the second spot?

OpenStudy (zachnovo15):

(-9,-3)

OpenStudy (ello1221):

thx i got some more questions though

OpenStudy (zachnovo15):

Alright. I can help.

OpenStudy (zachnovo15):

Just tag me in a problem or question you have and I'll answer it

OpenStudy (ello1221):

k

OpenStudy (ello1221):

how

OpenStudy (zachnovo15):

@Zachnovo15

OpenStudy (ello1221):

6x + 7 = 67

OpenStudy (zachnovo15):

X=10

OpenStudy (ello1221):

x over four + 1 = –6

OpenStudy (zachnovo15):

X=-28

OpenStudy (ello1221):

Riverview Farms sold a combined total of 396 watermelons and cantaloupes. they sold 126 more watermelons than cantaloupes. Use n to represent the number of cantaloupes. Which equation could be used to solve for n? (Hint: First, think about the expression you would use to represent the number of watermelons, and then add the number of cantaloupes.)

OpenStudy (ello1221):

u there?

OpenStudy (zachnovo15):

Yeah. Imy just trying to figure it out. Because it would be a multi step equation i think. Any answers you can show me?

OpenStudy (ello1221):

n + 126 = 396 n – 126 = 396 2n – 126 = 396 2n + 126 = 396 these are choices

OpenStudy (zachnovo15):

The last choice 2n+126=396

OpenStudy (ello1221):

thx Suppose you are driving to visit a friend in another state. You are driving 65 miles per hour. You must drive 520 miles total. If you have already driven 195 miles, how long will it take you to reach your destination? Use h to represent the number of hours it will take to reach your destination. Use the equation 65h + 195 = 520.

OpenStudy (zachnovo15):

5 hours

OpenStudy (ello1221):

Suzy scored at least 18 points. (1 point) K ≤ 18 K < 18 K ≥ 18 K > 18

OpenStudy (ello1221):

thx yuve helped so much so far

OpenStudy (zachnovo15):

The third answer. And you're welcome. Make sure you give a medal :D

OpenStudy (ello1221):

k how?

OpenStudy (zachnovo15):

Just hit the best response button

OpenStudy (ello1221):

k

OpenStudy (ello1221):

Write an inequality for the following situation: No more than 5 books are in your backpack. (1 point) x > 5 x ≤ 5 x ≥ 5 x < 5

OpenStudy (zachnovo15):

Second one

OpenStudy (ello1221):

Solve the inequality. y + 3 ≤ 5 (1 point) y ≥ 2 y ≤ 8 y ≤ 2 y ≥ 8

OpenStudy (zachnovo15):

3rd answer

OpenStudy (ello1221):

im still here

OpenStudy (ello1221):

Write an inequality for the sentence. Then solve the inequality. Fifteen less than a number is at most 6. (1 point) x + 15 ≤ 6; x ≤ –9 x – 15 ≤ 6; x ≤ 21 x – 15 ≥ 6; x ≥ 21 x – 15 ≥ 6; x ≥ 16

OpenStudy (zachnovo15):

3rd answer

OpenStudy (ello1221):

it wont let me copy paste

OpenStudy (ello1221):

The maximum weight allowed per car on The Twister carnival ride is 263 pounds. Your friend weighs 86 pounds. To be able to ride in a car together, how much can you weigh? Write and solve the inequality. (1 point) x + 177 ≤ 263; at most 86 pounds x – 86 ≤ 263; at most 177 pounds x + 86 ≤ 263; at most 177 pounds x + 86 ≤ 177; at most 263 pounds

OpenStudy (zachnovo15):

3rd answer

OpenStudy (ello1221):

plz explain for dis one Complete the function table. Then write a rule for the function. function table

OpenStudy (ello1221):

this goes wih question

OpenStudy (zachnovo15):

It goes in order. It goes down by one. -5, -4, -3,-2, -1 as the input decreases by one the output decreases by one

OpenStudy (ello1221):

Write an inequality to represent the situation: The temperature stayed above –15°.

OpenStudy (zachnovo15):

T>-15

OpenStudy (ello1221):

plz explain

OpenStudy (zachnovo15):

Ok well the temperature needs to be above negative 15. It can't equal it but Ithe can be higher. So t (temperature) is greater than -15

OpenStudy (zachnovo15):

Make sense?

OpenStudy (ello1221):

yep

OpenStudy (zachnovo15):

Ok good :D

OpenStudy (ello1221):

Gabriella went skiing. She paid $35 to rent skis and $15 an hour to ski. If she paid a total of $95, how many hours did she ski? Write an equation to represent the situation and solve it. plz explain alot on deez ones

OpenStudy (ello1221):

Solve the equation. c over four– 5 = 4 er solve dis one i got the other

OpenStudy (zachnovo15):

C=36

OpenStudy (ello1221):

explain plz

OpenStudy (zachnovo15):

You have to first add 5 to both sides. So the the 5 cancels out and you geto c over 4 equals 9. Next you have to multiply 4 to both side . Now you have c=36

OpenStudy (ello1221):

Solve the inequality and describe the solution set. y – 6 ≥ 12 this is the last one

OpenStudy (zachnovo15):

So you add 6 to both sides and you have y is greater than or equal to 18. That means that y can be 18 or greater and the inequality will still be right.